How many solutions are there to the equation below? 5(x+10)-25=5x+25
Answer by juliane1(5)

(5 x + 5 * 10) - 25 = 5x + 25
(5x + 50) - 25 = 5x + 25
5x + 50 - 25 = 5x + 25
5x + 25 = 5x + 25
0=0